What was found
DUE TO NOISE IN FLIGHT DECK TOMC SUGGESTED CHANGING RECIRCULATION AIR FILTERS. REPLACED RECIRCULATION AIR FILTERS. L AND R PACK FLOW CONTROL VALVES NEED TO BE REPLACED FOR TROUBLESHOOTING. REPLACED L AND R FLOW CONTROL VALVES. OPS AND LEAK CHECK GOOD. CHECKED UPPER DUCT REGISTER FOR BLOCKAGE. REGISTER GOOD. BEGAN TROUBLESHOOTING PER SER TIP 21-22-00 PG 901. DEC LEVEL IN FLT DECK WITH RECIRC FANS ONLY ARE WITHIN LIMITS. WITH PACKS ON DB LEVEL IS 77. REPLACED L AND R PACK FLOW CONTROL VALVES. LEVEL NOW AT 74 DB. REPLACED RECIRC FAN FILTERS, NO HELP. FLIGHT DECK SUPPLY DUCT FLOW AND VELOCITY ARE GOOD. FOUND CAPT EYELET FLOW LOW, CLEANED SCREEN. FLIGHT DECK DB IS NOW 72 DB. ALSO PACK HI FLOW TEST PASSED METHOD 1. RPL LT AND RT FLOW CONTROL VLVS, FILTERS, CLEANED OUT RESTRICTED EYEBALL IN FLIGHT DECK. LOWERED FLIGHT DECK DB FROM 77DB TO 72DB. A-C OK FOR FLIGHT.